OpenStudy (anonymous):

Write the equation of the line with a slope of 3/7 passing through the point (9, -2).

OpenStudy (anonymous):

Please HELP!

OpenStudy (anonymous):

use the point slope formula sub in your info and solve for y=mx+b

OpenStudy (anonymous):

Canyou show me how?

OpenStudy (anonymous):

the point slope formula is y-y1=x(x-x1)

OpenStudy (anonymous):

So can you do this one to show me as an example?

OpenStudy (anonymous):

y-(-2)=3/7(x-9) y+2=3/7(x-9) y+2=3/7x-27/7 y=3/7x-27/7-2 y=3/7x-13/7

OpenStudy (anonymous):

So can you do this one to show me as an example?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

You have a point (h,k) = (9,-2) You have a slope m = 3/7 Plug those into the equation: \[y-k = m(x-h)\] That is the equation for the line going through that point with that slope.